# Axion-like Dark Matter Search with Space-based Gravitational Wave Detectors

### v1

@@ I Introduction
The endeavor to unravel the nature of dark matter, one of the most profound and enduring mysteries in modern physics Zwicky (1933); Bertone and Tait (2018), has driven scientists to explore a wide range of theoretical candidates, ranging from ultra-light scalar particles to primordial black holes. Among these possibilities, axion-like dark matter (ALDM) has emerged as a particularly compelling candidate Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Peccei (2008); Marsh (2016); Adams et al. (2022); O’Hare (2024), distinguished by a unique set of properties that make it an attractive focus for both theoretical and experimental investigations Arvanitaki et al. (2010); Graham et al. (2015); Di Luzio et al. (2020). These particles are postulated to possess an extraordinarily low mass, exhibit exceedingly feeble interactions, and exist in large abundance.
@@
One of the most intriguing aspects of ALDM is its interaction with electromagnetic fields, which gives rise to distinctive signatures that are potentially detectable. A key phenomenon associated with ALDM is the birefringence effect Carroll et al. (1990); Carroll (1998); Harari and Sikivie (1992), which arises from the coupling between the ALDM field and photons. In the presence of an ALDM field, the phase velocities of left-handed and right-handed circularly polarized light differ while the light propagating through space. This difference causes linearly polarized light to undergo a rotation of its polarization plane as it propagates. This effect provides a distinctive and measurable signature of ALDM. By studying this birefringence effect, researchers can probe the strength of the ALDM-photon coupling, offering a powerful tool to test the existence of these elusive particles and constrain their intrinsic properties.
@@
While traditional astrophysical observations, such as cosmic microwave background (CMB) polarization studies Harari and Sikivie (1992); Fedderke et al. (2019); Fujita et al. (2021) and celestial polarimetry di Serego Alighieri et al. (2010); Fujita et al. (2019); Ivanov et al. (2019); Liu et al. (2020); Poddar and Mohanty (2020); Chen et al. (2022); Yuan et al. (2021); Liu et al. (2023), have been employed to search for birefringence effects induced by ALDM, their sensitivity is ultimately limited by systematic uncertainties and astrophysical foregrounds. To achieve greater precision, new experimental approaches leveraging high-precision interferometry have been proposed. In particular, optical cavities and laser interferometers DeRocco and Hook (2018); Obata et al. (2018); Liu et al. (2019); Nagano et al. (2019); Michimura et al. (2020); Heinze et al. (2024), which can measure tiny phase shifts with extraordinary accuracy, present promising avenues for detecting the subtle imprint of ALDM-induced birefringence.
@@
Gravitational wave (GW) detectors are well-developed interferometers for probing the phase signature induced by ALDM. The detection of GW has revolutionized astrophysics since the first observation by the LIGO-Virgo collaboration in 2015 Abbott et al. (2016), opening a novel observational window into the cosmos. Ground-based interferometers, such as LIGO Aasi et al. (2015), Virgo Acernese et al. (2015), and KAGRA Akutsu et al. (2019), now routinely detect high-frequency GWs ($10\,\mathrm{Hz}\sim 10\,\mathrm{kHz}$) originating from compact binary mergers, including black holes and neutron stars, while pulsar timing arrays McLaughlin (2013); Manchester et al. (2013); Kramer and Champion (2013); Manchester (2013) probe the nanohertz regime in search of stochastic signals Agazie et al. (2023); Reardon et al. (2023); Antoniadis et al. (2023); Xu et al. (2023) from supermassive black hole binaries. However, the millihertz frequency band—home to mergers of massive black hole binaries, extreme mass-ratio inspirals, and Galactic compact binaries—remains inaccessible to current ground-based or PTA efforts. This gap will be addressed by future space-based detectors, such as LISA Colpi et al. (2024), Taiji Luo et al. (2020), TianQin Luo et al. (2016), and Big-Bang Observer (BBO) Crowder and Cornish (2005), which utilize laser interferometry between spacecraft separated by millions of kilometers to achieve unprecedented phase sensitivity. These missions employ advanced technologies, including picometer-level metrology, drag-free spacecraft, and time-delay interferometry (TDI) Tinto and Dhurandhar (2021), to suppress noise sources, thereby enabling not only GW astronomy but also synergies with fundamental physics.
@@
Given the exceptional phase sensitivity of SGWDs, typically at the picometer level over million-kilometer baselines, these instruments hold significant potential for detecting the subtle effects induced by the ALDM. Lately, modifications to these SGWDs have been proposed to probe ALDM Gué et al. (2025); Yao et al. (2025). These studies suggest altering the polarization of the laser beam exchanged between the spacecraft (S/C) from linear to circular. With these modificaitons, the phase difference induced by ALDM would be integrated into the phase measurement of SGWDs. However, such modifications may introduce some issues like the susceptibility to roll rotations of the S/C Müller (2017) and the generation of backreflection LIS (2000).
@@
The paper is organized as follows. In Section II we review the ALDM-induced birefringence effect and its impact on the polarization state. Section III introduces our proposed modification to the optical bench design. Section IV presents a detailed sensitivity analysis of the ALDM-photon coupling. Section V discusses practical considerations. Finally, the conclusion is given in Section VI. Throughout this paper, all equations are expressed in Lorentz–Heaviside units.
II ALDM-induced Birefringence
In this section, we briefly introduce the birefringence effect induced by ALDM and its implications for space-based interferometry. The ALDM-photon interaction is decribed by the Lagrangian Carroll et al. (1990); Carroll and Field (1991); Harari and Sikivie (1992)
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{1}{2}\partial^{\mu}a\partial_{\mu}a-\frac{1}{2}m_{a}^{2}a^{2}-\frac{1}{4}g_{a\gamma}aF_{\mu\nu}\tilde{F}^{\mu\nu},$ (1)
@@ II ALDM-induced Birefringence
$\omega_{\pm}\simeq k\pm g_{a\gamma}\left(\frac{\partial a}{\partial t}+\nabla a\cdot\bm{\hat{k}}\right)=k\pm g_{a\gamma}n^{\mu}\nabla_{\mu}a,$ (2)
where $n^{\mu}$ denotes the null tangent vector to the photon path. This effect is known as ALDM-induced birefringence.
@@
The phase difference between the left and right circular polarization can be probed by SGWDs, constituting the central detection principle proposed in Gué et al. (2025); Yao et al. (2025). In their proposals, the linearly polarized light transmitted between S/C is substituted with circularly polarized light, and the phase shift induced by the birefringence effect is integrated into the interferometer signal. However, although the current design of SGWDs utilize linearly polarized laser links, the past designs of SGWDs favored circularly polarized light for the inter-satellite path Müller (2017), and both options were under consideration ESA (2011). It seems that the advantages of linear polarization were deemed more favorable after detailed trade-off studies. Using circular polarization requires converting from linear to circular via quarter-wave plates and then back to linear at the receiving end. Extra conversion steps are sensitive to relative rotations, coating phase shifts, and alignment errors between spacecraft. Circular polarization in phase sensitive ranging yields a susceptibility to roll rotations of the S/C Müller (2017). Besides, optical components, such as quarter-wave plates, are main sources of backreflection, and it is easier to design and to predict the behaviour of the antireflection coatings of the optics working only with linear polarisation LIS (2000). This motivates the exploration of alternative approaches that preserve the primary light path of SGWDs.
@@
[…]
@@
Neglecting the back-action term of light on the ALDM, the wave function of the ALDM filed, derived from the equations of motion arising from Eq. (1), can be written as
$a(t,\bm{x})=a_{0}\cos(\omega_{a}t-\bm{k}_{a}\cdot\bm{x}+\Phi),$ (4)
where $a_{0}$ is the field amplitude, related to the dark matter density $\rho_{\mathrm{DM}}=m_{a}^{2}a_{0}^{2}/2\approx 0.4\>\mathrm{GeV\cdot cm^{-3}}$ McMillan (2011), and $\Phi$ denotes an unknown phase. In the non-relativistic limit, the frequency $\omega_{a}$ and wave vector $\bm{k}_{a}$ are given by $\omega_{a}\simeq m_{a}(1+\bm{v}^{2}/2)$ and $\bm{k}_{a}\simeq m_{a}\bm{v}$, respectively, where $\bm{v}$ is the collective velocity of the ALDM with $v\sim 10^{-3}$. When the coherence length of the ALDM field, $\lambda_{a}=2\pi/k_{a}$, significantly exceeds the distance between S/C, denoted as $L$, Eq. (4) can be approximated as
$a(t)\approx\frac{\sqrt{2\rho_{\mathrm{DM}}}}{m_{a}}\cos(m_{a}t+\Phi).$ (5)
@@
$\Delta\phi_{a}=$ $\frac{1}{2}g_{a\gamma}\frac{\sqrt{2\rho_{\mathrm{DM}}}}{m_{a}}\left\{\cos(m_{a}t+\Phi)-\cos\left[m_{a}(t-L)+\Phi\right]\right\}$ (6)
@@
$\sim$ $-10^{-7}\left(\frac{g_{a\gamma}}{10^{-11}\,\mathrm{GeV}^{-1}}\right)\left(\frac{\rho_{\mathrm{DM}}}{0.4\,\mathrm{GeV}\cdot\mathrm{cm}^{-3}}\right)^{1/2}\left(\frac{10^{-16}\,\mathrm{eV}}{m_{a}}\right)$
×$\times\sin\left(\frac{m_{a}L}{2}\right)\sin\left(m_{a}t+\Phi-\frac{m_{a}L}{2}\right).$
@@
Without loss of generality, we suppose that the transmitting laser light from remote S/C is s-polarized with an amplitude $E_{\mathrm{TX}}$. The received light affected by the ALDM oscillation can be written as
$\bm{E}_{\mathrm{RX}}=\cos\Delta\phi_{a}E_{\mathrm{TX}}\hat{\bm{s}}+\sin\Delta\phi_{a}E_{\mathrm{TX}}\hat{\bm{p}},$ (7)
where $\hat{\bm{s}}$ and $\hat{\bm{p}}$ represent the unit vectors in the s-polarization and p-polarization directions, respectively. The presence of the ALDM field induces a small p-polarization component proportional to $\sin\Delta\phi_{a}$, while slightly reducing the s-polarization component proportional to $\cos\Delta\phi_{a}$.
From Eq. (7), the s-polarized component of the received light undergoes periodic modulation at $\mathcal{O}(\Delta\phi_{a}^{2})$ due to the ALDM field. However, this modulation lies below the sensitivity threshold of the current design of SGWDs. A possible approach to make the interferometer sensitive to the birefringence effect is to measure the amplitude of the p-polarized component of the received light, as proposed in studies such as Nagano et al. (2019). Nevertheless, direct detection of the p-polarized signal poses significant challenges in SGWDs, due to the extremely low power of the received light, which is on the order of $\sim 700\,\mathrm{pW}$. Our proposal circumvents this limitation by converting the polarization rotation into a differential phase measurement through the implementation of an auxiliary interferometer.
III Proposed Modification
Figure 1: Schematic of the modified optical bench design. The dotted box highlights the additional ALDM interferometer, phase shifter, and associated optics added to the current OB configuration. The orange beam represents the transmitted light, the green dotted path indicates the received s-polarized light, and the green waving path shows the received p-polarized light generated by ALDM-induced birefringence. PBS: polarizing beam splitter; BS: beam splitter; $\lambda/2$: half-wave plate; PAAM: point-ahead angle mechanism; A,B,C,D: photodector.
@@ III Proposed Modification
The main innovation of our approach lies in maintaining the established linear polarization scheme while adding a dedicated detection path for the ALDM-induced p-polarization component. This strategy preserves the advantages of the current SGWD design while enhancing sensitivity to new physics.
The proposed modification to the OB is illustrated in Fig. 1. The optical elements enclosed within the dotted line, including the ALDM interferometer, a phase shifter, and associated mirrors, are added to the OB, while the remaining components represent a simplified version of the current OB design for SGWDs. Polarizing beam splitters (PBS) are employed to separate the transmitted and received light. In the schematic, the transmitted light is depicted in orange, the received s-polarized light is represented by green dotted lines, and the received p-polarized light is illustrated with green wavy lines.
In our proposed design, the PBS directs the p-polarized component, which contains the ALDM signal, into the auxiliary interferometer path, while allowing the s-polarized component to propagate along the standard interferometer path. The auxiliary path includes a phase shifter that enables precise control of the relative phase between the two interferometer outputs, which is crucial for optimizing the detection sensitivity.
For the inter-spacecraft interferometer (ISI), the interference between a local transmitted beam and the received beam results in heterodyne signals at the photodiodes. In the presence of the ALDM field, the received amplitude is modulated by a factor of $\cos\Delta\phi_{a}$, while in the ALDM interferometer, the modulation appears as $\sin\Delta\phi_{a}$. With the auxiliary channel, the birefringence-induced amplitude modulation is converted into a phase shift $\phi_{\mathrm{ALDM}}$, which can be read out by combining the heterodyne outputs of the ISI and the ALDM interferometer appropriately. The phase shift signal is then compared with that from the ISI, $\phi_{\mathrm{ISI}}$, as follows
$\phi_{\mathrm{ALDM}}-\phi_{\mathrm{ISI}}\simeq\Delta\phi_{a},$ (8)
@@
In the ISI, a fraction of the local transmitted beam $\sqrt{2I_{0}}e^{i\omega_{\mathrm{T}}t}$ interferes with the light received from the remote S/C $p_{\mathrm{ISI}}\sqrt{2I_{0}}e^{i(\omega_{R}t+\phi)}$, where $I_{ref}=2I_{0}$ is the power of the local reference, $p_{\mathrm{ISI}}^{2}\sim 10^{-7}$ is the ratio of the received power to the local reference, $\omega_{\mathrm{T}}$ is the angular frequency of the transmitted beam, $\omega_{R}$ is the angular frequency of the received beam, and $\phi$ is the phase of the received wavefront. In the presence of the ALDM field, the power of the received light in the ISI is modulated by the birefringence effect, expressed as
$p_{\mathrm{ISI}}^{\prime}=\cos\Delta\phi_{a}p_{\mathrm{ISI}}.$ (9)
@@
$E_{A}=\sqrt{2I_{0}}e^{i\omega_{\mathrm{T}}t}\left[1+p_{A}^{\prime}e^{i(\Omega t+\phi+\pi/2)}+u_{A}e^{i(\psi+\pi/2)}\right],$ (10)
$E_{B}=\sqrt{2I_{0}}e^{i\omega_{\mathrm{T}}t}\left[e^{i\pi/2}+p_{B}^{\prime}e^{i(\Omega t+\phi)}+u_{B}e^{i\psi}\right],$ (11)
where the subscripts A and B label the diode ports, $\Omega=\omega_{R}-\omega_{\mathrm{T}}$ is the heterodyne frequency, $p_{\mathrm{ISI}}^{\prime 2}=p_{A}^{\prime 2}+p_{B}^{\prime 2}\ll 1$, the fraction $u_{\mathrm{ISI}}^{2}=u_{A}^{2}+u_{B}^{2}\ll 1$ and phase $\psi$ characterizes the field of the backscattered light. The heterodyne signals at the photodiode A and B are
$S_{A}(t)=2I_{0}\left[1+p_{A}^{\prime 2}+u_{A}^{2}-2u_{A}\sin\psi-2p_{A}^{\prime}\sin(\Omega t+\phi)+2p_{A}^{\prime}u_{A}\cos(\Omega t+\phi-\psi)\right],$ (12)
$S_{B}(t)=2I_{0}\left[1+p_{B}^{\prime 2}+u_{B}^{2}+2u_{B}\sin\psi+2p_{B}^{\prime}\sin(\Omega t+\phi)+2p_{B}^{\prime}u_{B}\cos(\Omega t+\phi-\psi)\right].$ (13)
@@
$S_{B}(t)-S_{A}(t)$ $=2I_{0}\left[p_{B}^{\prime 2}-p_{A}^{\prime 2}+u_{B}^{2}-u_{A}^{2}+2(u_{B}+u_{A})\sin\psi\right.$ (14)
@@
From the integrated output, the phase of the heterodyne-signal can be recovered as Sasso et al. (2019)
$\phi_{\mathrm{ISI}}\approx\phi+\pi/2+\tilde{u}\cos\psi-\tilde{u}^{2}\sin\psi\cos\psi,$ (16)
with
$\tilde{u}_{\mathrm{ISI}}=\frac{p_{A}^{\prime}u_{A}-p_{B}^{\prime}u_{B}}{p_{A}^{\prime}+p_{B}^{\prime}}.$ (17)
Similarly, for the ALDM interferometer, the signal at the photodiode C and D is given by
$S_{D}-S_{C}$ $=2I_{1}\left[p_{D}^{\prime 2}-p_{C}^{\prime 2}+u_{D}^{2}-u_{C}^{2}+2(u_{D}+u_{C})\sin\tilde{\psi}\right.$ (18)
$\left.+2(p_{D}^{\prime}+p_{C}^{\prime})\sin(\Omega t+\phi+\Delta\phi)+2(p_{D}^{\prime}u_{D}-p_{C}^{\prime}u_{C})\cos(\Omega t+\phi+\Delta\phi-\tilde{\psi})\right]$

### v2

@@ 1 Introduction
The endeavor to unravel the nature of dark matter, one of the most profound and enduring mysteries in modern physics [Zwicky:1933gu, Bertone:2018krk], has driven scientists to explore a wide range of theoretical candidates, ranging from ultra-light scalar particles to primordial black holes. Among these possibilities, axion-like dark matter (ALDM) has emerged as a particularly compelling candidate [Preskill:1982cy, Abbott:1982af, Dine:1982ah, Peccei:2006as, Marsh:2015xka, Adams:2022pbo, OHare:2024nmr], distinguished by a unique set of properties that make it an attractive focus for both theoretical and experimental investigations [Arvanitaki:2009fg, Graham:2015ouw, DiLuzio:2020wdo]. These particles are postulated to possess an extraordinarily low mass, exhibit exceedingly feeble interactions, and exist in large abundance.
@@
One of the most intriguing aspects of ALDM is its interaction with electromagnetic fields, which gives rise to distinctive signatures that are potentially detectable. A key phenomenon associated with ALDM is the birefringence effect [Carroll:1989vb, Carroll:1998zi, Harari:1992ea], which arises from the coupling between the ALDM field and photons. In the presence of an ALDM field, the phase velocities of left-handed and right-handed circularly polarized light differ while the light propagates through space. This difference causes linearly polarized light to undergo a rotation of its polarization plane as it propagates. This effect provides a distinctive and measurable signature of ALDM. By studying this birefringence effect, researchers can probe the strength of the ALDM-photon coupling, offering a powerful tool to test the existence of these elusive particles and constrain their intrinsic properties.
@@
While traditional astrophysical observations, such as cosmic microwave background (CMB) polarization studies [Harari:1992ea, Fedderke:2019ajk, Fujita:2020aqt] and celestial polarimetry [diSeregoAlighieri:2010ejm, Fujita:2018zaj, Ivanov:2018byi, Liu:2019brz, Poddar:2020qft, Chen:2021lvo, Yuan:2020xui, Liu:2021zlt, Gan:2023swl], have been employed to search for birefringence effects induced by ALDM, their sensitivity is ultimately limited by systematic uncertainties and astrophysical foregrounds. To achieve greater precision, new experimental approaches leveraging high-precision interferometry have been proposed. In particular, optical cavities and laser interferometers [DeRocco:2018jwe, Obata:2018vvr, Liu:2018icu, Nagano:2019rbw, Michimura:2019qxr, Heinze:2023nfb], which can measure tiny phase shifts with extraordinary accuracy, present promising avenues for detecting the subtle imprint of ALDM-induced birefringence.
@@
Gravitational wave (GW) detectors are well-developed interferometers for probing the phase signature induced by ALDM. The detection of GW has revolutionized astrophysics since the first observation by the LIGO-Virgo collaboration in 2015 [LIGOScientific:2016aoc], opening a novel observational window into the cosmos. Ground-based interferometers, such as LIGO [LIGOScientific:2014pky], Virgo [VIRGO:2014yos], and KAGRA [KAGRA:2018plz], now routinely detect high-frequency GWs ($10\,\mathrm{Hz}\sim 10\,\mathrm{kHz}$) originating from compact binary mergers, including black holes and neutron stars, while pulsar timing arrays [McLaughlin:2013ira, Manchester:2012za, Kramer:2013kea, Manchester:2013ndt] probe the nanohertz regime in search of stochastic signals [NANOGrav:2023gor, Reardon:2023gzh, EPTA:2023fyk, Xu:2023wog] from supermassive black hole binaries. However, the millihertz frequency band—home to mergers of massive black hole binaries, extreme mass-ratio inspirals, and Galactic compact binaries—remains inaccessible to current ground-based or PTA efforts. This gap will be addressed by future space-based detectors, such as LISA [LISA:2024hlh], Taiji [Luo:2019zal], TianQin [TianQin:2015yph], and Big-Bang Observer (BBO) [Crowder:2005nr], which utilize laser interferometry between spacecraft separated by millions of kilometers to achieve unprecedented phase sensitivity. These missions employ advanced technologies, including picometer-level metrology, drag-free spacecraft, and time-delay interferometry (TDI) [Tinto:2020fcc], to suppress noise sources, thereby enabling not only GW astronomy but also synergies with fundamental physics.
@@
Given the exceptional phase sensitivity of SGWDs, typically at the picometer level over million-kilometer baselines, these instruments hold significant potential for detecting the subtle effects induced by the ALDM. Lately, modifications to these SGWDs have been proposed to probe ALDM [Gue:2024txz, Yao:2024hap]. These studies suggest altering the polarization of the laser beam exchanged between the spacecraft (S/C) from linear to circular. With these modifications, the phase difference induced by ALDM would be integrated into the phase measurement of SGWDs. However, such modifications may introduce some issues like the susceptibility to roll rotations of the S/C [muller2017design] and the generation of backreflection [LISA2000ftr].
@@
The paper is organized as follows. In Section 2 we review the ALDM-induced birefringence effect and its impact on the polarization state. Section 3 introduces our proposed modification to the optical bench design. Section 4 presents a detailed sensitivity analysis of the ALDM-photon coupling. Section 5 discusses practical considerations. Finally, the conclusion is given in Section 6. Throughout this paper, all equations are expressed in Lorentz–Heaviside units.
2 ALDM-induced Birefringence
In this section, we briefly introduce the birefringence effect induced by ALDM and its implications for space-based interferometry. The ALDM-photon interaction is described by the Lagrangian [Carroll:1989vb, Carroll:1991zs, Harari:1992ea]
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{1}{2}\partial^{\mu}a\partial_{\mu}a-\frac{1}{2}m_{a}^{2}a^{2}-\frac{1}{4}g_{a\gamma}aF_{\mu\nu}\tilde{F}^{\mu\nu},$ (2.1)
@@ 2 ALDM-induced Birefringence
$\omega_{\pm}\simeq k\pm g_{a\gamma}\left(\frac{\partial a}{\partial t}+\nabla a\cdot\bm{\hat{k}}\right)=k\pm g_{a\gamma}n^{\mu}\nabla_{\mu}a,$ (2.2)
where $n^{\mu}$ denotes the null tangent vector to the photon path. [^1] The simplified expression found in other works is a special case of Eq. (2.2), valid when the spatial gradient term $\nabla a\cdot\hat{\mathbf{k}}$ is negligible. This assumption is valid since we are considering the coherence length of the ALDM field significantly exceeds the distance between S/C. This effect is known as ALDM-induced birefringence.
@@
The phase difference between the left and right circular polarization can be probed by SGWDs, constituting the central detection principle proposed in [Gue:2024txz, Yao:2024hap]. In their proposals, the linearly polarized light transmitted between S/C is substituted with circularly polarized light, and the phase shift induced by the birefringence effect is integrated into the interferometer signal. However, although the current design of SGWDs utilizes linearly polarized laser links, the past designs of SGWDs favored circularly polarized light for the inter-satellite path [muller2017design], and both options were under consideration [ESA2011LISAYellowBook]. It seems that the advantages of linear polarization were deemed more favorable after detailed trade-off studies. Using circular polarization requires converting from linear to circular via quarter-wave plates and then back to linear at the receiving end. Extra conversion steps are sensitive to relative rotations, coating phase shifts, and alignment errors between spacecraft. Circular polarization in phase sensitive ranging yields a susceptibility to roll rotations of the S/C [muller2017design]. Besides, optical components, such as quarter-wave plates, are main sources of backreflection, and it is easier to design and to predict the behaviour of the antireflection coatings of the optics working only with linear polarization [LISA2000ftr]. This motivates the exploration of alternative approaches that preserve the primary light path of SGWDs.
@@
$\Delta\phi_{a}=$ $\frac{1}{2}\int_{t_{\mathrm{emit}}}^{t_{\mathrm{obs}}}(\omega_{+}-\omega_{-})dt$ (2.3)
@@
Neglecting the back-action term of light on the ALDM, the wave function of the ALDM field, derived from the equations of motion arising from Eq. (2.1), can be written as
$a(t,\bm{x})=a_{0}\cos(\omega_{a}t-\bm{k}_{a}\cdot\bm{x}+\Phi),$ (2.4)
where $a_{0}$ is the field amplitude, related to the dark matter density $\rho_{\mathrm{DM}}=m_{a}^{2}a_{0}^{2}/2\approx 0.4\>\mathrm{GeV\cdot cm^{-3}}$ [McMillan:2011wd], and $\Phi$ denotes an unknown phase. In the non-relativistic limit, the frequency $\omega_{a}$ and wave vector $\bm{k}_{a}$ are given by $\omega_{a}\simeq m_{a}(1+\bm{v}^{2}/2)$ and $\bm{k}_{a}\simeq m_{a}\bm{v}$, respectively, where $\bm{v}$ is the collective velocity of the ALDM with $v\sim 10^{-3}$. When the coherence length of the ALDM field, $\lambda_{a}=2\pi/k_{a}$, significantly exceeds the distance between S/C, denoted as $L$, Eq. (2.4) can be approximated as
$a(t)\approx\frac{\sqrt{2\rho_{\mathrm{DM}}}}{m_{a}}\cos(m_{a}t+\Phi).$ (2.5)
@@
$\Delta\phi_{a}=$ $\frac{1}{2}g_{a\gamma}\frac{\sqrt{2\rho_{\mathrm{DM}}}}{m_{a}}\left\{\cos(m_{a}t+\Phi)-\cos\left[m_{a}(t-L)+\Phi\right]\right\}$ (2.6)
@@
[…]
$\times\sin\left(\frac{m_{a}L}{2}\right)\sin\left(m_{a}t+\Phi-\frac{m_{a}L}{2}\right).$
@@
Without loss of generality, we suppose that the polarization angle between the transmitting laser light from remote S/C and the optical bench reference basis is $\theta=\theta_{0}+\Delta\phi_{a}$, where $\theta_{0}$ is the misalignment polarization angle without ALDM. The received light affected by the ALDM oscillation can be written as
$\bm{E}_{\mathrm{RX}}\propto\cos\theta\hat{\bm{s}}+\sin\theta\hat{\bm{p}}$. (2.7)
Here $\hat{\bm{s}}$ and $\hat{\bm{p}}$ represent the unit vectors in the s-polarization and p-polarization directions, respectively.
In general setup of SGWDs, the angle $\theta_{0}$ is calibrated to be 0, then
$\bm{E}_{\mathrm{RX}}\propto\cos\Delta\phi_{a}\hat{\bm{s}}+\sin\Delta\phi_{a}\hat{\bm{p}}$. (2.8)
The presence of the ALDM field induces a small p-polarization component proportional to $\sin\Delta\phi_{a}$, while slightly reducing the s-polarization component proportional to $\cos\Delta\phi_{a}$. From Eq. (2.8), the s-polarized component of the received light undergoes periodic modulation at $\mathcal{O}(\Delta\phi_{a}^{2})$ due to the ALDM field. However, this modulation lies below the sensitivity threshold of the current design of SGWDs. A possible approach to make the interferometer sensitive to the birefringence effect is to measure the amplitude of the p-polarized component of the received light, as proposed in studies such as [Nagano:2019rbw]. Nevertheless, direct detection of the p-polarized signal poses significant challenges in SGWDs, due to the extremely low power of the received light, which is on the order of $\sim 700\,\mathrm{pW}$. Our proposal circumvents this limitation by converting the polarization rotation into a differential phase measurement through the implementation of an auxiliary interferometer.
3 Proposed Modification
Figure 1: Schematic of the modified optical bench design. The dotted box highlights the additional auxiliary interferometer, phase shifter, and associated optics added to the current OB configuration. The orange beam represents the transmitted light, the green dotted path indicates the received s-polarized light, and the green waving path shows the received p-polarized light. The arrows show the beam direction of propagation, whereas the double-tipped arrows and dots represent the in-plane and out-of-plane polarizations. PBS: polarizing beam splitter; BS: beam splitter; $\lambda/2$: half-wave plate; PAAM: point-ahead angle mechanism; A, B, C, D: photodetector.
@@ 3 Proposed Modification
The main innovation of our approach lies in maintaining the established linear polarization scheme while adding a dedicated detection path for the p-polarization component. This strategy preserves the advantages of the current SGWD design while enhancing sensitivity to new physics.
The proposed modification to the OB is illustrated in Fig. 1. The optical elements enclosed within the dotted line, including the auxiliary interferometer, a phase shifter, and associated mirrors, are added to the OB, while the remaining components represent a simplified version of the current OB design for SGWDs. Polarizing beam splitters (PBS) are employed to separate the transmitted and received light. In the schematic, the transmitted light is depicted in orange, the received s-polarized light is represented by green dotted lines, and the received p-polarized light is illustrated with green wavy lines. The proposed optical design is primarily illustrative, and a full engineering design of the optical bench is beyond the scope of this study.
In our proposed design, the received p-polarized component is converted to s-polarized and then directed by the PBS into the auxiliary interferometer path, while the s-polarized component propagates along the standard interferometer path. The auxiliary path includes a phase shifter that enables precise control of the relative phase between the two interferometer outputs, which is crucial for optimizing the detection sensitivity. The TX dump suppresses the back-scattered light from the received p-polarized light.
For SGWDs, the interference between a local transmitted beam and the received beam results in heterodyne signals at the photodiodes. In the presence of the ALDM field, the polarization direction of the received light is $\theta=\theta_{0}+\Delta\phi_{a}$. In the inter-spacecraft interferometer (ISI), the received amplitude is modulated by a factor of $\cos\theta$, while in the auxiliary interferometer (AUXI), the modulation appears as $\sin\theta$. With the auxiliary channel, the amplitude modulation due to the polarization angle can be converted into phase signals $\phi_{\mathrm{AUXI}}$ and $\phi_{\mathrm{ISI}}$. The phase $\phi_{\mathrm{AUXI}}$ is read out by combining the heterodyne outputs of the ISI and the auxiliary interferometer appropriately while $\phi_{\mathrm{ISI}}$ is the phase signal from the ISI. Comparing these two signals, one can recover the polarization angle
$\theta\simeq\phi_{\mathrm{AUXI}}-\phi_{\mathrm{ISI}},$ (3.1)
@@
In the ISI, a fraction of the local transmitted beam $\sqrt{2I_{0}}e^{i\omega_{\mathrm{T}}t}$ interferes with the light received from the remote S/C $p_{\mathrm{ISI}}\sqrt{2I_{0}}e^{i(\omega_{R}t+\phi)}$, where $I_{ref}=2I_{0}$ is the power of the local reference, $p_{\mathrm{ISI}}^{2}\sim 10^{-7}$ is defined as the ratio of the received optical power to the local reference beam when the incoming light is in a pure s-polarized state, $\omega_{\mathrm{T}}$ is the angular frequency of the transmitted beam, $\omega_{R}$ is the angular frequency of the received beam, and $\phi$ is the phase of the received wavefront. In the presence of the ALDM field, the power of the received light in the ISI is modulated by the birefringence effect, expressed as
$p_{\mathrm{ISI}}^{\prime}=\cos\theta p_{\mathrm{ISI}}.$ (3.2)
@@
$E_{A}=\sqrt{2I_{0}}e^{i\omega_{\mathrm{T}}t}\left[1+p_{A}^{\prime}e^{i(\Omega t+\phi+\pi/2)}+u_{A}e^{i(\psi+\pi/2)}\right],$ (3.3)
$E_{B}=\sqrt{2I_{0}}e^{i\omega_{\mathrm{T}}t}\left[e^{i\pi/2}+p_{B}^{\prime}e^{i(\Omega t+\phi)}+u_{B}e^{i\psi}\right],$ (3.4)
where the subscripts A and B label the diode ports, $\Omega=\omega_{R}-\omega_{\mathrm{T}}$ is the heterodyne frequency, $p_{\mathrm{ISI}}^{\prime 2}=p_{A}^{\prime 2}+p_{B}^{\prime 2}\ll 1$, the fraction $u_{\mathrm{ISI}}^{2}=u_{A}^{2}+u_{B}^{2}\ll 1$ and phase $\psi$ are the fractional power and phase of the backscattered light field. The heterodyne signals at the photodiodes A and B are
$S_{A}(t)=2I_{0}\left[1+p_{A}^{\prime 2}+u_{A}^{2}-2u_{A}\sin\psi-2p_{A}^{\prime}\sin(\Omega t+\phi)+2p_{A}^{\prime}u_{A}\cos(\Omega t+\phi-\psi)\right],$ (3.5)
$S_{B}(t)=2I_{0}\left[1+p_{B}^{\prime 2}+u_{B}^{2}+2u_{B}\sin\psi+2p_{B}^{\prime}\sin(\Omega t+\phi)+2p_{B}^{\prime}u_{B}\cos(\Omega t+\phi-\psi)\right].$ (3.6)
@@
$S_{B}(t)-S_{A}(t)$ $=2I_{0}\left[p_{B}^{\prime 2}-p_{A}^{\prime 2}+u_{B}^{2}-u_{A}^{2}+2(u_{B}+u_{A})\sin\psi\right$. (3.7)
@@
$\phi_{\mathrm{ISI}}=\arg\left[\int_{0}^{2n\pi/\Omega}\left[S_{B}(t)-S_{A}(t)\right]e^{-i\Omega t}dt\right],$ (3.8)
From the integrated output, the phase of the heterodyne-signal can be recovered as [Sasso:2018bau]
$\phi_{\mathrm{ISI}}\approx\phi+\pi/2+\tilde{u}\cos\psi-\tilde{u}^{2}\sin\psi\cos\psi,$ (3.9)
with
$\tilde{u}_{\mathrm{ISI}}=\frac{p_{A}^{\prime}u_{A}-p_{B}^{\prime}u_{B}}{p_{A}^{\prime}+p_{B}^{\prime}}.$ (3.10)
Similarly, for the AUXI, the signal at the photodiode C and D is given by
$S_{D}-S_{C}$ $=2I_{1}\left[p_{D}^{\prime 2}-p_{C}^{\prime 2}+u_{D}^{2}-u_{C}^{2}+2(u_{D}+u_{C})\sin\tilde{\psi}\right$. (3.11)